Terry has 7 more quarters than dimes and has a total of $6.80. The number of quarters and dimes is 38. How many quarters and dimes does terry have?
Thanks for answering!:-)

Answer by fractalier(6550)

Let q = # of quarters. Thus their value is 25q.
Let d = # of dimes. Thus their value is 10d.
From the facts in the problem, we have
q + d = 38
25q + 10d = 680 ($6.80 expressed in cents)
We also have a third equation, not common in these problems (q = d + 7) which incidentally is impossible...
Let us solve the first for d and substitute into the second equation...
d = 38 - q
25q + 10(38 - q) = 680
25q + 380 - 10q = 680
15q = 300
q = 20 and then
d = 18